Status Report

NASA Space Station On-Orbit Status 28 Feb 2004

By SpaceRef Editor
February 28, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 28 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.   Flight Control to Crew: “One of the best and purest joys is having a rest after labor.  (Immanuel Kant)”

Saturday,- nominally a day of sleeping long.  After the “nap”, which had been inserted in the crew’s workday, ended this morning at 3:00am EST, Foale and Kaleri performed a number of tasks before having their dinner at ~10:30am this morning and beginning their “regular” sleep cycle at 12:00pm, for a wakeup early tomorrow at 1:00am (which gets them back on the nominal sleep cycle).

Before sleep time, the crew complete EVA-9 closeout operations by packing up and stowing the retrieved folded Japanese MPAC/SEED container for return to Earth.

Sasha started another regeneration cycle on absorbent bed #1 of the BMP harmful impurities unit in the SM, leaving channel 2 in Purify mode.

Mike took the regular ppCO2 (carbon dioxide partial pressure) readings in the SM and Lab with the CDMK (carbon dioxide monitoring kit).   [These weekly readings in both modules, using the same CDMK unit, are now being conducted for the next few weeks to help the ground better assess if IMV (intermodular ventilation) flow between the U.S. and Russian segments has become degraded.]

The crew had the option of working out on CEVIS bike and TVIS treadmill, and Mike then downloaded the accumulated data files from the exercise equipment to the MEC via memory card and RED log entries, for downlink on OCA comm.  The CDR also performed the periodic (weekly) transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, and then deleted them on the HRM.  

TsUP/Moscow confirmed that the following objectives were accomplished during the abbreviated EVA-9 on 2/26: (1) replacement of an SKK removable materials sample cassette from DC-1, (2) retrieval of Japanese MPAC/SEEDS panel #2 and relocation of MPAC/SEEDS panel #3, (3) installation/activation of the ESA “Matryoshka” experiment, (4) removal of two Velcro (Aramid) straps from the WA2 ham radio antenna, and (5) replacement of two SKK removable materials sample cassettes on the Service Module (SM). 
Not completed were: (1) retrieval/relocation of Automated Transfer Vehicle (ATV) Laser Retro Reflectors (LRR) for ESA, and (2) retrieval/installation of “Kromka” thruster contamination sampling tray; two additional tasks involving the “Platan-M” payload experiment and imagery of external structures, which were to be performed only if the other tasks were completed early, were also not accomplished.  All incoimplete tasks can be included in the two EVAs planned for Increment 9.  

When SM panel power was turned off during EVA preps, the PCS (portable computer system) laptop in the SM was also deactivated, which was in violation of a flight rule and the EVA-9 Minimum Equipment List.   [The flight rule requires a minimum of two active PCS’s attached to core data busses for Caution & Warning support.  After coordinated evaluation/assessment of the situation by MCC-H specialists, including the IMMT Chair, the violation was deemed acceptable, and EVA ops continued as planned.  The PCS was successfully reactivated following the spacewalk.]

During EVA-9 preparations, the PromISS-3 crystal growth experiment in the MSG (Microgravity Science Glovebox) experienced an unexpected shutdown.  MSFC/POIC (Payload Operations & Integration Center) recovered PromISS to its nominal configuration with a ground commanded reboot (see below).

Following the pressure equalization between the U.S. and Russian segments after EVA-9, ground controllers at MCC-H noted a 0.5-psi pressure differential between the CL (Crew Lock) and the EL (Equipment Lock) in the U.S. “Quest” Airlock, which was below the 0.7-psi flight rule limit.  The CDR was instructed to relieve the excess pressure using the MEV (manual equalization valve) on the A/L’s IV hatch prior to crew sleep.

Weekly Science Update (Expedition Eight — 17th):

GASMAP:   Next GASMAP operations will take place in mid-March..

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound:   Planned.

Hand Posture Analyzer (HPA:  Looking forward to future operations.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Looking forward to future operations, after the PromISS operations.
Space Acceleration Measurement System (SAMS):   SAMS captured Orlan EVA data on 2/26; recovered from ICU laptop lock-up.
Microgravity Acceleration Measurement System (MAMS):   MAMS captured Orlan EVA data on 2/26.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):  Experimental conditions are optimal and the optical microscope is showing very stable fringe patterns.  A new table was uploaded on 2/24 that modified the video recording rate in such a way that tape #12 can be used for 6 days.  The hardware is now recording one wheel revolution for every four.  Although scheduled for 2/24 at 8:20pm EST, the tape exchange from #11 to #12 occurred at 2/25 (7:28am), but minimal science loss is expected, since no more than two revolutions of the wheel were not recorded.  This is insignificant compared to the approximately 420 revolutions to be recorded in all.  On 2/26, an unexpected reset of the experiment was experienced, during which the hardware reverted to a warm-up state after 26 days of nominal ops.  After this reset, telemetry immediately showed that everything looked 100% nominal.  The science loss due to the reset appears to be very limited.  Tape #12 will most likely run out by the afternoon of 3/2.  PromISS will automatically go to “Wait Tape” mode if the tape runs out.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Renal Stone (RS):   Last in-flight data collection session will occur in April.  Thanks to the crew for their continued support with taking their daily pills at dinner time.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   Pre-Increment requirements have been completed.  A second BBT (Beacon & Beacon Tester) session will be scheduled in US Lab.

Foot/Ground Reaction Forces during Space Flight (FOOT):   The Foot Data will be downlinked a second time next week.  As soon as the PI has seen the data, Mike Foale will be provided with any updates and address his questions after the last Ops session.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Looking forward to the next set of FDI Tissue Culture runs.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Nothing new. 

Earth Science Toward Exploration Research (ESTER):  Looking forward to the upcoming sessions.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   CGBA continues to run nominally at 20C.

Miscible Fluids in Microgravity (MFMG):    Nothing new.

Educational Payload Operations (EPO):  The education demonstrations continue to be excellent.  The video will be used in a variety of ways.  The ground is looking forward to upcoming activities.

Crew Earth Observations (CEO):   A striking ISS/CEO image of the village of Argudan on the north slopes of the Caucasus Mountains will be published on Earth Observatory this weekend as low light and snowcover enhance the distinctive terrain and land use patterns of this region.  Also published this past week in the Dynamic Events section of Earth Observatory were the crew’s dramatic views of two storms in southwestern Asia on 2/15.  The first was a sandstorm sweeping over Qatar on the Persian Gulf and the second was a large dust storm over the Kerman Desert of southeastern Iran.  Careful review of ISS/CEO imagery determined that the crew recently acquired views of three of CEO city targets that meet requirements: Bogota, Columbia; Phoenix, AZ; and a particularly outstanding Seattle, WA.  The crew also made a fine effort in documenting aerosol events taking advantage of their varying look angle, lighting conditions, and lens settings.  While the ground has seen improvement in image focus overall, the crew is to keep working on their long-lens views.  Camera times remain consistent and correct.

ISS Orbit  (as of this noon, 12:42pm EST [= epoch]):

  • Mean altitude — 365.7 km
  • Apogee — 369.7 km
  • Perigee — 361.6 km
  • Period — 91.86 min.
  • Inclination (to Equator) —  51.629 deg
  • Eccentricity — 0.0006004
  • Orbits per 24-hr. day — 15.68
  • Mean altitude loss last 24 hours — 135 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 30114

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.